The square of the speed of an object undergoing a uniform acceleration a is some function of a and the displacement s, according to the expression

Velocity squared = k*acceleration power m*Speed power n

What dimensions should k have in order for the expression to be dimensionally consistent ?
1
Expert's answer
2020-02-06T09:35:12-0500

Velocity squared is measured in "\frac{m^2}{s^2}" , acceleration - in "\frac{m}{s^2}" , so acceleration to power m - in "\frac{m^m}{s^{2m}}" , speed - in "\frac{m}{s}" , so speed to power n - in "\frac{m^n}{s^n}" . So, speaking in dimensions:

"\frac{m^2}{s^2}=k\cdot \frac{m^m}{s^{2m}}\cdot \frac{m^n}{s^n}"

"k=\frac{m^2}{s^2}:\frac{m^{m+n}}{s^{2m+n}}"


Answer. The dimension of k is

"k=\frac{m^{2-m-n}}{s^{2-2m-n}}"